OpenStudy (anonymous):

A problem based on mean value theorem. Please ping me. 1) Suppose that f'(x) >= 0 for all x > a. Prove that f(x) >= f(a) for all x >= a?

OpenStudy (anonymous):

if x = a, we have satisfied the requirement (obviously) if x > a then by the MVT \(\exists\) c, \(a<c<x\) s.t.\[f'\left( c \right)=\frac{ f \left( x \right)-f \left( a \right) }{ x-a }>0\] let me knwo if you need the rest

OpenStudy (anonymous):

assuming f is continuous on the interval [a, x]

OpenStudy (anonymous):

Can u send me the full problem in detail please

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

since f'(x) \(\ge\) 0, \(\forall\) x > a => f(x) - f(a) \(\ge\) 0 (because x - a > 0) => f(x) \(\ge\) f(a)